Given the following assumptions, develop a cost estimate (Labor only) that integrates all of the cost elements into a final bid price to the customer. What is the total cost for the project? How long will the project take to complete?
Function points in the project: 232
Software engineers assigned to this team: 5
Function point productivity per software engineer: 5 per month
Workdays per typical month: 22
Productive hours per typical workday: 8
Gross hourly wage rate per software engineer (does not include fringe benefits): $50 per hour
Overhead (fringe benefit, other direct overhead) rate 35%
G&A rate: 11.5%
Proposed Profit: 10%
Answer #1

Monthly productivity of team = 5x5 =25

Time taken by the team to complete entre project = 232 /25 =9.28 month

Number of workdays in these months = 22*9.8 =204.16

Number of work hours = 8x204.16 =1633.28 =1634 hours

Gross hourly wages for team = 1634 x 50x5 =408500

Overhead and G&A = (35+11.5)% of wages = 189952.5

Total cost = 598452.5

Total bid = 1.1 x 598452.5 =658297.75

=658298

Project will tak 9.28 months